Please help me with this multiple choice question:

Which quadratic equation has the roots 3+i and 3-i?

a) x^2 + 6x-10 = 0 b) x^2 + 6x + 8= 0 c) x^2 - 6x+10 = 0 d) x^2- 6x-8 =0
thank you!

Found 2 solutions by edjones, solver91311:
Answer by edjones(8007)

roots 3+i and 3-i
x=3+i, x=3-i
x-3+i=0
x-3-i=0
(x-(3+i))(x-(3-i))=0
x^2-6x+10=0 c)
.
Ed

Answer by solver91311(24713)


If a quadratic equation has roots and , then and are factors of the quadratic polynomial, so:



are the factors of the polynomial that you are looking for. Just multiply it out and collect like terms to find the appropriate quadratic polynomial. Hint: Consider 3 + i and 3 - i as single numbers and use FOIL. The product of those two numbers will be the difference of two squares. Also remember that

John